2. An initially stationary cart rolls down an incline with a measured acceleration of 2.7 m/s. How long does it take the cart to roll 85 m down the incline? 13 2019 Motion with Constant Acceleration
Answer #1

Given,

a = 2.7 m/s^2

We know that

S = ut + 1/2 at^2

here, s = 85 m ;

85 = 0 + 1/2 x 2.7 t^2

t = sqrt (2 x 85/2.7) = 7.93 m/s^2

Hence, t = 7.93 s